# thanks: This work is based on the author’s M.Sc. dissertation (Ref.[1]), but further correcting and expanding upon it

### v1

@@ III Hessian Diagonalization and Sector Classification
where $\mathbf{1}$ denotes a 2-by-2 identity matrix block. If $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\neq 0$ but $M_{S}\gg m$, we find
$\mathcal{H}=\begin{pmatrix}\left[\left(\Pi_{\mu}+\eta S_{\mu}\right)^{2}+2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right]\mathbf{1}&0\\ 0&\left[\left(\Pi_{\mu}-\eta S_{\mu}\right)^{2}+2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right]\mathbf{1}\end{pmatrix}$ (38)
which exhibits an unstable vacuum for one chirality and conversely for the other. In time-like configurations of the axial vector, the Hessian exhibits these same signs of instability, even for a very massive axial vector, leading to the expression:
$\lambda^{(\pm)}_{\pm}\approx\left(\pm\right)2\sqrt{\eta m\left(-\eta mS^{2}\pm\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right)}$ (39)
This implies that the eigenvalues invariably contain at least two pure imaginary components, provided the condition $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\geq\eta mS^{2}$ holds. If attainable, this scenario indicates a chirality-asymmetric stability, where parity is violated in a manner that allows one chirality sector to remain “safe”  [^1] This safety can be put into question if higher loops, interactions, and backreactions are included according to more complicated scenarios. while the opposite sector invariably decays into particles. This interplay results in a chirality-asymmetric pair creation, characterized by a net production of fermions with a preferred chirality. Should any process propel $\eta mS^{2}$ beyond the remaining term, the system will become entirely unstable, precipitating rapid vacuum decay.
Significantly, this sector also demonstrates a kinematical symmetry reduction to the $SIM(2)$ (or $ISO(2)$ if the norm of $S_{\mu}$ is fixed) subgroup of the Lorentz group. This assertion is underscored by the structure of Eq. (37), which bears resemblance to that found in Very Special Relativity [60]. Similar symmetry breaking is also observed in the preceding subspaces, though not with the same level of explicitness and specificity. Consequently, models constructed within this subspace have the potential to exhibit a rich phenomenology, including the intriguing possibility of imparting mass to neutrinos without necessitating a Seesaw mechanism [61], provided that careful considerations are made during the model’s development [*[WhiletheHessianinquestioniswell-behaved, itmaynotencompassalldesiredrepresentations, includingprojectiveones, asdiscussedby][.]Marcondes, *Marcondes2].
@@
Apart from the previously discussed subspaces, further stable regions may exist elsewhere in parameter space, contingent upon careful considerations and specific choices informed by the structure of Eq. (28). Given that the combined regions of MT- and NonAnom-types encompass a significant portion of the parameter space, the remaining regions are either special cases within/transitional between these two categories—such as those involving the vanishing of additional invariants or potentially exhibiting complex eigenvalue structures—or intrinsically unstable for any parameter values. Noteworthy regions that fall under the category of exhibiting long eigenvalue expressions are strictly governed by conditions such as $S^{\mu}=0$, $\mathcal{V}=0$, $\left(V^{\mu\nu}S_{\nu}\right)^{2}=0$, or any combinations thereof, without imposing additional constraints on other invariants.
@@
$\mathcal{V}\geq 2m^{2}\eta^{2}S^{2}\pm$
$\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$
@@
$\pm 4\eta\sqrt{\left(\Pi_{\mu}S^{\mu}\right)^{2}\mathcal{V}+m^{2}\left(\tilde{V}^{\mu\nu}S_{\nu}\right)}$
@@
$\eta\left(\Pi_{\mu}S^{\mu}\right)^{2}\geq\eta m^{2}S^{2}\pm m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$
Possibly Chirality-Asymmetric-Unstable
@@
$\mathcal{V}\geq 2\eta^{2}m^{2}\eta mS^{2}\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$ (40)
holds true. Importantly, this subspace reduces to the EQED-type for massless axial-vector theories where $S^{2}=0$. The TRMW-type remains fully stable for Electroaxial theories involving massless pseudo-vectors, presenting intriguing and physically consistent characteristics for other massive axial-vector theories. Nonetheless, it permits the potential for chirality-asymmetric pair creation when a system with massive $S_{\mu}$ evolves into it from another subspace. Conversely, massive pseudo-vector theories within the NonAnom-type (including TRMW-type) may not be considered fundamental—primarily due to the presence of $\Pi_{\mu}S^{\mu}$ within $\mathcal{H}_{\mathrm{R}}$—and the condition $M_{S}\gg m$ must be satisfied along with constraints on $\mathcal{V}$ and/or $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$ to ensure the existence of at least one stable chirality sector. Further stability constraints may be required for different scenarios where $\mathrm{S}_{0}\left[A,S\right]$ could affect the consistency requirements or in cases where spacetime curvature is taken into account, or when both vector fields are quantized [64, *Nonpertrenorm, *StabilitydeRham].
@@ IV The Exact One-Loop Effective Action
Since both fields are massless, we can utilize Fock-Schwinger’s gauge for each field, such that $V^{(R/L)}_{\mu}=-\left(eF_{\mu\nu}\pm\eta S_{\mu\nu}\right)x^{\nu}/2\pm\bar{S}_{\mu}$. If $F_{\mu\nu}$ and $S_{\mu\nu}$ form a regular pencil [*[See, e.g., ][forfurtherdetails.]gantmacher], it is possible to simultaneously transform them into Jordan’s form through an appropriate choice of coordinates. By applying boosts and rotations, we can choose a reference frame where each vector potential can be parameterized as
@@
Furthermore, since $\mathcal{V}=\mathcal{W}=0$, whereby we find $E^{(H)}_{V}=\pm B^{(H)}_{V}$ when $\mathcal{F}^{(H)}_{V}=\mathcal{G}^{(H)}_{V}=0$, that is, they form plane waves. Alternatively, if we allow the fields to take on complex values, we find $B^{(H)}_{V}=iE^{(H)}_{V}=\sqrt{\mathcal{F}^{(H)}_{V}}$. Thus, we can infer that Eq. (37) reduces to
$\mathcal{H}_{\mathrm{H}}=\left[p_{t}-E^{(H)}_{V}x\right]^{2}-p^{2}_{x}-\left[p_{y}-B^{(H)}_{V}z\right]^{2}-p_{z}^{2}-X_{(H)}$ (44)
where $X_{(R)}=-2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\mu}\right|}\,\sigma_{z}=\mathcal{X}_{+}\sigma_{z}$ and $X_{(L)}=-2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\mu}\right|}\,\sigma_{z}=\mathcal{X}_{-}\sigma_{z}$. Next, by decomposing the $n$-th eigenstate of $\mathcal{H}$ into its chirality components $\left|\psi_{n}\right>=\left|\psi_{R,n}\right>+\left|\psi_{L,n}\right$>, we can express the trace as $\mathrm{Tr}e^{-i\tau\mathcal{H}}=\mathrm{Tr}_{\mathrm{R}}e^{-i\tau\mathcal{H}_{\mathrm{R}}}+\mathrm{Tr}_{\mathrm{L}}e^{-i\tau\mathcal{H}_{\mathrm{L}}}$ and employ momentum operators as
@@
$\mathcal{H}_{\mathrm{H}}=-\left[p^{2}_{x}-\left(E^{(H)}_{V}\right)^{2}x^{2}\right]-\left[p^{2}_{z}+\left(B^{(H)}_{V}\right)^{2}z^{2}\right]-X_{(H)}$ (46)
@@
where $X=\mathrm{diag}\left\{\mathcal{X}_{+}\sigma_{z},\mathcal{X}_{-}\sigma_{z}\right\}$, $X_{\pm}=\left|\lambda^{\left(\pm\right)}_{\pm}\right|$, and $\sigma_{z}=\mathrm{diag}\left\{1,-1\right\}$ represent the Pauli matrix in the $z$-direction. The next step involves decomposing $S_{\mu}$ into its transverse and longitudinal modes, followed by adopting an approximate Fock-Schwinger gauge for its transverse component within the domain defined by the quasi-static assumptions, specifically $S_{\mu}=S_{\mu\nu}x^{\nu}+\bar{S}_{\mu}$, satisfying $S^{2}=\left(S_{\mu\nu}x^{\nu}\right)^{2}+\bar{S}^{2}\approx\bar{S}^{2}$. We can extract $\bar{S}^{2}$ from the Hessian, leading us to define the effective mass as
@@
Consequently, it is feasible to parameterize $A_{\mu}=\left(-E_{A}x,0,-B_{A}z,0\right)$ and $S_{\mu}^{\perp}=\left(-E_{S}x,0,-B_{S}z,0\right)$ or interchangeably shift $z\to x$ for either vector. We can then apply the momentum operators as to shift
@@
To facilitate our analysis, we assume integration within a Euclidean box of size $L$ and calculate the number of eigenstates contained within: (i) for the first class, each chirality sector may possess a different density of eigenstates, approximately given by $\sim 4\pi^{2}/\left(\mathcal{G}^{(H)}_{V}L^{2}\right)$, whereas (ii) for the second class, at energies below the fermion mass, the factors multiplying the momenta in Eq. (52) will have a minimal impact on the dispersion relation, leading to a density of eigenstates approximately $\sim 4\pi^{2}/\left(\mathcal{G}L^{2}\right)$, where $\mathcal{G}=EB$. Consequently, we derive
@@
which, upon returning to the expression $L^{4}\to\int d^{4}x$ and Wick rotating $\tau\to i\tau$, ultimately yields
$\overline{\Gamma}_{c}\left[A,S\right]=\frac{1}{16\pi^{2}}\sum_{(H)}\int d^{4}x\ \frac{d\tau}{\tau}\left[\frac{e^{-M_{c}^{2}\tau}\mathcal{G}^{(H)}_{I}\cos{\left(\tau\mathcal{X}_{H}\right)}}{\sin{\left(\tau E^{(H)}_{I}\right)}\sinh{\left(\tau B^{(H)}_{I}\right)}}-\frac{e^{-m^{2}\tau}}{\tau^{2}}\right]$ (55)
Here, we denote the mass term of class $c$ as $M_{c}$, specifically $M_{1}=m$ or $M_{2}=M$. More explicitly, in the first class, the whole quantum correction vanishes for plane wave solutions, while it takes the form
@@
$\left.+\frac{\left(E^{(L)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(L)}_{V}\right)}}-\frac{2}{\tau^{2}}\right]$
@@
for the second class. For this, we have defined $\bar{\mathcal{X}}_{\pm}=\sqrt{2\mathcal{F}\pm 2i\mathcal{G}}=B\pm iE$, where $\mathcal{F}=\frac{1}{2}(B^{2}-E^{2})$. To employ a minimal subtraction scheme, we initially expand both components around $\tau\to 0$, resulting in
@@
$\left.+\frac{\left(E^{(L)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(L)}_{V}\right)}}-\frac{2}{\tau^{2}}-\frac{1}{3}\left[\left(E^{(R)}_{V}\right)^{2}+\left(E^{(L)}_{V}\right)^{2}\right]\right\}$
@@
Before elaborating on this result, it is important to recognize that due to $S_{\mu}$ being a pseudo-vector, the mixed term in $\mathcal{F}_{V}^{(H)}$ ($\mathcal{G}_{V}^{(H)}$) does (not) violate CP invariance, contrasting with the corresponding term found in $\mathcal{V}$ ($\mathcal{W}$). Thus, additional constraints, such as $S^{\mu\nu}F_{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$, may be necessary for model consistency [69]. If this constraint holds, we deduce that $\mathcal{F}^{(R)}_{V}=\mathcal{F}^{(L)}_{V}=\mathcal{V}$ and $\mathcal{G}^{(R)}_{V}=\mathcal{G}^{(L)}_{V}=\mathcal{W}$, which leads to the conclusion that Eq.(60) vanishes since it occurs when $\mathcal{V}=\mathcal{W}=0$, implying there are no surviving one-loop quantum corrections to the classical action. In other words, an effective action resembling the EH form for the TRMW-type subspace is exclusively feasible within CP-violating theories.
For the remainder of this section, we will operate within a MT-type subspace to elucidate these results, assuming $F_{\mu\nu}S^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=\left(\tilde{V}^{\mu\nu}S_{\nu}\right)^{2}=0$ for simplicity, where $X_{+}=\sqrt{2\mathcal{V}-4m^{2}\eta^{2}S^{2}+i\mathcal{W}}$. Thus, Eq. (61) transforms into
@@
Thus, we conclude that Eq. (67) is consistent with the minimal subtraction term of Eq. (62) in the limit of constant fields since $\Delta_{M}=\int d\tau/\tau e^{-M^{2}\tau}=\Delta-\ln{\left(1-\eta^{2}S^{2}/m^{2}\right)}\approx\Delta$. It is noteworthy that the third-order term also conforms to expectations, despite a naive assumption that it might not be present if the analysis is kept merely at the expansion of Eq. (62), as will be elaborated in the following section.

### v2

@@ III Hessian Diagonalization and Sector Classification
where $\mathbf{1}$ denotes a 2-by-2 identity matrix block. If $|\widetilde{V}^{\mu\nu}S_{\nu}|\neq 0$ but $M_{S}\gg m$, we find again Eq. 33. In time-like configurations of the axial vector, the Hessian exhibits these same signs of instability of Eq. 32, even for a very massive axial vector. This implies that the eigenvalues invariably contain at least two pure imaginary components, provided the condition $|\widetilde{V}^{\mu\nu}S_{\nu}|\geq\eta mS^{2}$ holds. If attainable, these scenarios indicate a chirality-asymmetric stability, where parity is violated in a manner that allows one chirality sector to remain “safe”  [^1] This safety can be put into question if higher loops, interactions, and backreactions are included according to more complicated scenarios. while the opposite sector invariably decays into particles. This interplay results in a chirality-asymmetric pair creation, characterized by a net production of fermions with a preferred chirality. Should any process propel $\eta mS^{2}$ beyond the remaining term, the system will become entirely unstable, precipitating rapid vacuum decay.
Significantly, this sector and the similar region within MT, where Eq. (33) is valid, also demonstrates a kinematical symmetry reduction to the SIM(2)SIM(2)—or $ISO(2)$ if the norm of $S_{\mu}$ is fixed—subgroup of the Lorentz group. This assertion is underscored by the structure of Eqs. (33) and (42), which bears resemblance to that found in Very Special Relativity [67]. Similar symmetry breaking is also observed in the preceding subspaces, though not with the same level of explicitness and specificity. Consequently, models constructed within this subspace have the potential to exhibit a rich phenomenology, including the intriguing possibility of imparting mass to neutrinos without necessitating a Seesaw mechanism [68], provided that careful considerations are made during the model’s development [*[WhiletheHessianinquestioniswell-behaved, itmaynotencompassalldesiredrepresentations, includingprojectiveones, asdiscussedby][.]Marcondes, *Marcondes2].
@@
Apart from the previously discussed subspaces, further stable regions may exist elsewhere in parameter space, contingent upon careful considerations and specific choices informed by the structure of Eq. (28). Given that the combined regions of MT- and NonAnom-types encompass a significant portion of the parameter space (even intersecting each other), the remaining regions are either special cases within/transitional between these two categories—such as those involving the vanishing of additional invariants or potentially exhibiting intricate eigenvalue structures—or intrinsically unstable for any parameter values. Noteworthy regions that fall under the category of exhibiting long eigenvalue expressions are strictly governed by conditions such as $S^{\mu}=0$, $\mathcal{V}=0$, $(V^{\mu\nu}S_{\nu})^{2}=0$, or any combinations thereof, without imposing additional constraints on other invariants.
@@
$2m^{2}\eta^{2}S^{2}\leq\mathcal{V}+$
$\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$
Possibly Chirality-Asymmetric
@@
$\pm 4\eta\sqrt{\left(\Pi_{\mu}S^{\mu}\right)^{2}\mathcal{V}+m^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)}$
@@
$\eta\left(\Pi_{\mu}S^{\mu}\right)^{2}+m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\geq\eta m^{2}S^{2}$
Possibly Chirality-Asymmetric
@@
$\mathcal{V}+\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}\geq 2\eta^{2}m^{2}\eta mS^{2}$ (43)
holds true. Importantly, this subspace reduces to the EQED-type for massless axial-vector theories where $S^{2}=0$ and share common regions with both NonAnom- and TRMW-type subspaces. The latter remains fully stable for Electroaxial theories involving massless pseudo-vectors, presenting intriguing and physically consistent characteristics for other massive axial-vector theories. Nonetheless, both TRMW- and MT-type permits the potential for chirality-asymmetric pair creation when a system with massive $S_{\mu}$ evolves into it from another subspace. Conversely, massive pseudo-vector theories within the NonAnom-type (including TRMW-type) may not be considered fundamental—primarily due to the presence of $\Pi_{\mu}S^{\mu}$ within $\mathcal{H}_{\mathrm{R}}$—and the condition $M_{S}\gg m$ must be satisfied along with constraints on $\mathcal{V}$ and/or $|\widetilde{V}^{\mu\nu}S_{\nu}|$ to ensure the existence of at least one stable chirality sector. Further stability constraints may be required for different scenarios where $\mathrm{S}_{0}[A,S]$ could affect the consistency requirements or in cases where spacetime curvature is taken into account, or when both vector fields are quantized [71, *Nonpertrenorm, *StabilitydeRham].
@@ IV The Exact One-Loop Effective Action
Since both fields are massless, we can utilize Fock-Schwinger’s gauge for each field, such that $V^{(R/L)}_{\mu}=-(eF_{\mu\nu}\pm\eta S_{\mu\nu})x^{\nu}/2\pm\bar{S}_{\mu}$. If $F_{\mu\nu}$ and $S_{\mu\nu}$ form a regular pencil [*[See, e.g., ][forfurtherdetails.]gantmacher], it is possible to simultaneously transform them into Jordan’s form through an appropriate choice of coordinates. By applying boosts and rotations, we can choose a reference frame where each vector potential can be parameterized as
@@
Furthermore, since $\mathcal{V}=\mathcal{W}=0$, whereby we find $E^{(H)}_{V}=\pm B^{(H)}_{V}$ when $\mathcal{F}^{(H)}_{V}=\mathcal{G}^{(H)}_{V}=0$, that is, they form plane waves. Alternatively, if the fields are allowed to take on complex values, we find $B^{(H)}_{V}=iE^{(H)}_{V}=\sqrt{\mathcal{F}^{(H)}_{V}}$. Thus, we can infer that Eq. (42) reduces to
$\mathcal{H}_{\mathrm{H}}=\left[p_{t}-E^{(H)}_{V}x\right]^{2}-p^{2}_{x}-\left[p_{y}-B^{(H)}_{V}z\right]^{2}-p_{z}^{2}+X_{(H)}$ (47)
where $X_{(R)}=2\sqrt{\eta m|\widetilde{V}^{\mu\nu}S_{\mu}|}\,\sigma_{z}=\mathcal{X}_{+}\sigma_{z}$ and $X_{(L)}=2i\sqrt{\eta m|\widetilde{V}^{\mu\nu}S_{\mu}|}\,\sigma_{z}=\mathcal{X}_{-}\sigma_{z}$. Next, by decomposing the $n$-th eigenstate of $\mathcal{H}$ into its chirality components $\left|\psi_{n}\right>=\left|\psi_{R,n}\right>+\left|\psi_{L,n}\right$>, we can express the trace as $\mathrm{Tr}e^{-i\tau\mathcal{H}}=\mathrm{Tr}_{\mathrm{R}}e^{-i\tau\mathcal{H}_{\mathrm{R}}}+\mathrm{Tr}_{\mathrm{L}}e^{-i\tau\mathcal{H}_{\mathrm{L}}}$ and employ momentum operators as
@@
$\mathcal{H}_{\mathrm{H}}=-\left[p^{2}_{x}-\left(E^{(H)}_{V}\right)^{2}x^{2}\right]-\left[p^{2}_{z}+\left(B^{(H)}_{V}\right)^{2}z^{2}\right]+X_{(H)}$ (49)
@@
where $X=\mathrm{diag}\{\mathcal{X}_{+}\sigma_{z},\mathcal{X}_{-}\sigma_{z}\}$ and $X_{\pm}=|\lambda^{\left(\pm\right)}_{\pm}|$. The next step involves decomposing $S_{\mu}$ into its transverse and longitudinal modes, followed by adopting an approximate Fock-Schwinger gauge for its transverse component within the domain defined by the quasi-static assumptions, specifically $S_{\mu}=S_{\mu\nu}x^{\nu}+\bar{S}_{\mu}$, satisfying $S^{2}=(S_{\mu\nu}x^{\nu})^{2}+\bar{S}^{2}\approx\bar{S}^{2}$. We can extract $\bar{S}^{2}$ from the Hessian, leading us to define the effective mass as
@@
Consequently, it is feasible to parameterize $A_{\mu}=(-E_{A}x,0,-B_{A}z,0)$ and $S_{\mu}^{\perp}=(-E_{S}x,0,-B_{S}z,0)$ or interchangeably shift $z\to x$ for either vector. We can then apply the momentum operators as to shift
@@
To facilitate our analysis, we assume integration within a Euclidean box of size $L$ and calculate the number of eigenstates contained within: (i) for the first class, each chirality sector may possess a different density of eigenstates, approximately given by $\sim 4\pi^{2}/(\mathcal{G}^{(H)}_{V}L^{2})$, whereas (ii) for the second class, at energies below the fermion mass, the factors multiplying the momenta in Eq. (55) will have a minimal impact on the dispersion relation, leading to a density of eigenstates approximately $\sim 4\pi^{2}/(\mathcal{G}L^{2})$, where $\mathcal{G}=EB$. Consequently, we derive
@@
which, upon returning to the expression $L^{4}\to\int d^{4}x$ and Wick rotating $\tau\to-i\tau$, ultimately yields
$\overline{\Gamma}_{c}\left[A,S\right]=\frac{1}{16\pi^{2}}\sum_{(H)}\int d^{4}x\ \frac{d\tau}{\tau}\left[\frac{e^{-M_{c}^{2}\tau}\mathcal{G}^{(H)}_{I}\cosh{\left(\tau\mathcal{X}_{H}\right)}}{\sin{\left(\tau E^{(H)}_{I}\right)}\sinh{\left(\tau B^{(H)}_{I}\right)}}-\frac{e^{-m^{2}\tau}}{\tau^{2}}\right]$ (58)
Here, we denote the mass term of class $c$ as $M_{c}$, specifically $M_{1}=m$ or $M_{2}=M$.
More explicitly, in the first class, the whole quantum correction vanishes for plane wave solutions, while it takes the form
@@
$\left.+\frac{\left(E^{(R)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(R)}_{V}\right)}}-\frac{2}{\tau^{2}}\right]$
@@
for the second class. For this, we have defined $\bar{\mathcal{X}}_{\pm}=\sqrt{2\mathcal{F}\pm 2i\mathcal{G}}=B\pm iE$, where $\mathcal{F}=1/2\,(B^{2}-E^{2})$. To employ a minimal subtraction scheme, we initially expand both components around $\tau\to 0$, resulting in
@@
$\left.+\frac{\left(E^{(R)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(R)}_{V}\right)}}-\frac{2}{\tau^{2}}-\frac{1}{3}\left[\left(E^{(R)}_{V}\right)^{2}+\left(E^{(L)}_{V}\right)^{2}\right]\right\}$
@@
Before elaborating on this result, it is important to recognize that due to $S_{\mu}$ being a pseudo-vector, the mixed term in $\mathcal{F}_{V}^{(H)}$ ($\mathcal{G}_{V}^{(H)}$) does (not) violate CP invariance, contrasting with the corresponding term found in $\mathcal{V}$ ($\mathcal{W}$). Thus, additional constraints, such as $S^{\mu\nu}F_{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$, may be necessary for model consistency [76]. If this constraint holds, we deduce that $\mathcal{F}^{(R)}_{V}=\mathcal{F}^{(L)}_{V}=\mathcal{V}$ and $\mathcal{G}^{(R)}_{V}=\mathcal{G}^{(L)}_{V}=\mathcal{W}$, which leads to the conclusion that Eq.(63) vanishes since it occurs when $\mathcal{V}=\mathcal{W}=0$, implying there are no surviving one-loop quantum corrections to the classical action. In other words, an effective action resembling the EH form for the TRMW-type subspace is exclusively feasible within CP-violating theories. If CP violation is permitted, then Eq. (33) can be simplified to a form akin to Eq. (63), yielding $E_{V}^{(R)}=E_{V}^{(L)}=E$. However, it is advisable to maintain the contributions from each sector distinct to accurately capture the chirality-asymmetric instability inherent in the particle creation process.
For the remainder of this section, we will operate within a MT-type subspace to elucidate these results, assuming $F_{\mu\nu}S^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=(\widetilde{V}^{\mu\nu}S_{\nu})^{2}=0$ for simplicity, where $X_{+}=\sqrt{2\mathcal{V}-4m^{2}\eta^{2}S^{2}+i\mathcal{W}}$. Thus, Eq. (64) transforms into
@@
Thus, we conclude that Eq. (70) is consistent with the minimal subtraction term of Eq. (65) in the limit of constant fields since $\Delta_{M}=\int d\tau\,e^{-M^{2}\tau}/\tau=\Delta-\ln{(1-\eta^{2}S^{2}/m^{2})}\approx\Delta$. It is noteworthy that the third-order term also conforms to expectations, despite a naive assumption that it might not be present if the analysis is kept merely at the expansion of Eq. (65), as will be elaborated in the following section.